OpenStudy (anonymous):

PLEASEEEE HELP I've been stuck on this problem MEDAL TO THE BEST ANSWER :) Indicate the equation of the line that is the perpendicular bisector of the segment with endpoints (4, 1) and (2, -5).

OpenStudy (aravindg):

First find 2 things. 1.slope of line joining these 2 points 2. The midpoint.

OpenStudy (anonymous):

So would the midpoint be (3,-2)

OpenStudy (aravindg):

Yep and slope?

OpenStudy (anonymous):

I don't know how to find the slope

OpenStudy (aravindg):

Slope=y difference/x difference

OpenStudy (anonymous):

2.3?

OpenStudy (aravindg):

Try again?

OpenStudy (anonymous):

I think im doing it wrong

OpenStudy (anonymous):

slope = (y2-y1)/(x2-x1)

OpenStudy (aravindg):

Lets take an example (5,2) and (3,7) Slope=2-7/5-3=-5/2

OpenStudy (anonymous):

-6 , -2 ?

OpenStudy (aravindg):

-6/-2=?

OpenStudy (anonymous):

3

OpenStudy (aravindg):

Good so slope of a perpendicular line will be -1/3

OpenStudy (anonymous):

But I have to put the answer like this

OpenStudy (aravindg):

Now you have point (3,-2) and slope -1/3 . Use slope point form and get ur line equation.

OpenStudy (anonymous):

y=(-1/3)-1?

OpenStudy (aravindg):

I think you missed out a term on left.

OpenStudy (anonymous):

the answer that said it's wrong because it would not be able to fit in the open spaces of the problem

OpenStudy (anonymous):

x+1?